OpenStudy (anonymous):

Solve for x: Logbase 8 (X+1)=2/3

OpenStudy (anonymous):

Change it to exponential : 8^2/3=(x+1) solve for x .

OpenStudy (anonymous):

\[\huge \log_{8}(x+1)=\frac{2}{3}\] Using the logarithm laws: \[\huge \log_{a}b=x\] The above expression can be written as: \[\huge a^x=b\] Using that, we cna write your logarithm equation as: \[\huge 8^{\frac{2}{3}}=x+1\] \[\huge \sqrt[3]{8^2}=x+1\] \[\huge \sqrt[3]{64}=x+1\] \[\huge 4=x+1\] Solve for x now.